# Channel-Imposed Fusion: A Simple yet Effective Method for Medical Time Series Classification

### v1

@@ 5.1 Results of Subject-Dependent
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| ADFTD (3-Classes) Reported | Autoformer Wu et al. (2021) | 87.831.62 | 87.631.66 | 87.221.97 | 87.381.79 | 96.590.88 | 93.821.64 |
| Crossformer Zhang and Yan (2022) | 89.351.32 | 89.001.44 | 88.791.37 | 88.881.40 | 97.520.58 | 95.451.03 | |
| FEDformer Zhou et al. (2022) | 77.632.37 | 76.762.17 | 76.682.48 | 76.602.46 | 91.671.34 | 84.942.11 | |
| Informer Zhou et al. (2021) | 90.930.90 | 90.740.71 | 90.501.14 | 90.600.94 | 98.190.27 | 96.510.49 | |
| iTransformer Liu et al. (2024b) | 64.900.25 | 62.530.27 | 62.210.26 | 62.250.33 | 81.520.29 | 68.870.49 | |
| MTST Zhang et al. (2024) | 65.080.69 | 63.850.80 | 62.710.64 | 63.030.58 | 81.360.56 | 69.340.89 | |
| Nonformer Liu et al. (2022) | 96.120.47 | 95.940.56 | 95.990.38 | 95.960.47 | 99.590.09 | 99.080.16 | |
| PatchTST Nie et al. (2023) | 66.260.40 | 65.080.41 | 64.970.51 | 64.950.42 | 83.070.45 | 71.700.61 | |
| Reformer Kitaev et al. (2019) | 91.511.75 | 91.151.79 | 91.651.56 | 91.141.83 | 98.850.35 | 97.880.60 | |
| Transformer Vaswani et al. (2017) | 97.000.43 | 96.870.53 | 96.860.36 | 96.860.44 | 99.750.04 | 99.420.07 | |
| Medformer Wang et al. (2024) | 97.620.34 | 97.530.33 | 97.480.40 | 97.500.36 | 99.830.05 | 99.620.12 | |
| MedGNN Fan et al. (2025) | 98.420.04 | 98.310.02 | 98.290.05 | 98.300.12 | 99.930.11 | - | |
| Medformer + CIF | 98.870.26 | 98.770.27 | 98.860.27 | 98.810.27 | 99.960.01 | 99.920.03 | |
@@
To further validate our proposed method, we reproduced 12 baseline approaches. Table 1 presents a comparison of the reported experimental results for these methods, while Table 5 in Appendix E provides detailed results from our reproduction of the 12 methods. The experimental results demonstrate that by incorporating our proposed CIF method into the Medformer model, we not only achieve significantly better performance than the original methods, but also outperform the current state-of-the-art (SOTA) approach, MedGNN. These findings strongly confirm the effectiveness and superiority of our method.
@@ 5.2 Results of Subject-Independent
In the Subject-Independent setup, the training, validation, and test sets are split based on subjects. All subjects and their corresponding samples are distributed into these three sets according to a predetermined ratio or subject IDs. Samples from the same subject exclusively appear in one of the three sets. This setup simulates real-world MedTS-based disease diagnosis, aiming to train a model on subjects with known labels and then test it on unseen subjects to determine if they have a specific disease. All five datasets are evaluated under this setup.
Table 2 presents the results reported by various methods in the subject-independent setting, while Table 6 shows the results of our reproduction of these methods. Our method achieves the highest scores in six metrics across four out of the five datasets. For the PTB-XL dataset, compared to the results reported by other methods, our approach achieves the highest AUROC and AUPRC score and ranks second in Accuracy. When compared to the results of our reproduction, our method ranks first in Accuracy, Precision, AUROC, and AUPRC, and second in Recall. Additionally, it is worth noting that in the subject-independent setup, the F1 score of ADFTD is 55.42%, which is significantly lower than the 99.59% achieved in the subject-dependent setup. This comparison highlights the challenges of the subject-independent setup, which better simulates real-world scenarios.
@@
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC |
| APAVA (2-Classes) Reported | Autoformer Wu et al. (2021) | 68.641.82 | 68.482.10 | 68.772.27 | 68.061.94 | 75.943.61 | 74.384.05 |
| Crossformer Zhang and Yan (2022) | 73.771.95 | 79.294.36 | 68.861.70 | 68.931.85 | 72.393.33 | 72.053.65 | |
| FEDformer Zhou et al. (2022) | 74.942.15 | 74.591.50 | 73.563.55 | 73.513.39 | 83.721.97 | 82.942.37 | |
| Informer Zhou et al. (2021) | 73.114.40 | 75.176.06 | 69.174.56 | 69.475.06 | 70.464.91 | 70.755.27 | |
| iTransformer Liu et al. (2024b) | 74.551.66 | 74.772.10 | 71.761.72 | 72.301.79 | 85.591.55 | 84.391.57 | |
| MTST Zhang et al. (2024) | 71.141.59 | 79.300.97 | 65.272.28 | 64.013.16 | 68.872.34 | 71.061.60 | |
| Nonformer Liu et al. (2022) | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |
| PatchTST Nie et al. (2023) | 67.031.65 | 78.761.28 | 59.912.02 | 55.973.10 | 65.650.28 | 67.990.76 | |
| Reformer Kitaev et al. (2019) | 78.702.00 | 82.503.95 | 75.001.61 | 75.931.82 | 73.941.40 | 76.041.14 | |
| Transformer Vaswani et al. (2017) | 76.304.72 | 77.645.95 | 73.095.01 | 73.755.38 | 72.506.60 | 73.237.60 | |
| Medformer Wang et al. (2024) | 78.740.64 | 81.110.84 | 75.400.66 | 76.310.71 | 83.200.91 | 83.660.92 | |
| MedGNN Fan et al. (2025) | 82.600.35 | 87.700.22 | 78.930.09 | 80.250.16 | 85.930.26 | - | |
| HM-BiTCN + CIF | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 | |
| TDBrain (2-Classes) Reported | Autoformer Wu et al. (2021) | 87.333.79 | 88.063.56 | 87.333.79 | 87.263.84 | 93.812.26 | 93.322.42 |
| Crossformer Zhang and Yan (2022) | 81.562.19 | 81.972.25 | 81.562.19 | 81.502.20 | 91.201.78 | 91.511.71 | |
| FEDformer Zhou et al. (2022) | 78.131.98 | 78.521.91 | 78.131.98 | 78.042.01 | 86.561.86 | 86.481.99 | |
| Informer Zhou et al. (2021) | 89.022.50 | 89.432.14 | 89.022.50 | 88.982.54 | 96.640.68 | 96.750.63 | |
| iTransformer Liu et al. (2024b) | 74.671.06 | 74.711.06 | 74.671.06 | 74.651.06 | 83.371.14 | 83.731.27 | |
| MTST Zhang et al. (2024) | 76.963.76 | 77.243.59 | 76.963.76 | 76.883.83 | 85.274.46 | 82.815.64 | |
| Nonformer Liu et al. (2022) | 87.882.48 | 88.861.84 | 87.882.48 | 87.782.56 | 97.050.68 | 96.990.68 | |
| PatchTST Nie et al. (2023) | 79.253.79 | 79.604.09 | 79.253.79 | 79.203.77 | 87.954.96 | 86.366.67 | |
| Reformer Kitaev et al. (2019) | 87.922.01 | 88.641.40 | 87.922.01 | 87.852.08 | 96.300.54 | 96.400.45 | |
| Transformer Vaswani et al. (2017) | 87.171.67 | 87.991.68 | 87.171.67 | 87.101.68 | 96.280.92 | 96.340.81 | |
| Medformer Wang et al. (2024) | 89.620.81 | 89.680.78 | 89.620.81 | 89.620.81 | 96.410.35 | 96.510.33 | |
| MedGNN Fan et al. (2025) | 91.040.09 | 91.150.12 | 91.040.20 | 91.040.08 | 96.740.04 | - | |
| HM-BiTCN + CIF | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 | |
| ADFTD (3-Classes) Reported | Autoformer Wu et al. (2021) | 45.251.48 | 43.671.94 | 42.962.03 | 42.591.85 | 61.021.82 | 43.102.30 |
| Crossformer Zhang and Yan (2022) | 50.452.31 | 45.571.63 | 45.881.82 | 45.501.70 | 66.452.03 | 48.332.05 | |
| FEDformer Zhou et al. (2022) | 46.300.59 | 46.050.76 | 44.221.38 | 43.911.37 | 62.621.75 | 46.111.44 | |
| Informer Zhou et al. (2021) | 48.451.96 | 46.541.68 | 46.061.84 | 45.741.38 | 65.871.27 | 47.601.30 | |
| iTransformer Liu et al. (2024b) | 52.601.59 | 46.791.27 | 47.281.29 | 46.791.13 | 67.261.16 | 49.531.21 | |
| MTST Zhang et al. (2024) | 45.602.03 | 44.701.33 | 45.051.30 | 44.311.74 | 62.500.81 | 45.160.85 | |
| Nonformer Liu et al. (2022) | 49.951.05 | 47.710.97 | 47.461.50 | 46.961.35 | 66.231.37 | 47.331.78 | |
| PatchTST Nie et al. (2023) | 44.370.95 | 42.401.13 | 42.061.48 | 41.971.37 | 60.081.50 | 42.491.79 | |
| Reformer Kitaev et al. (2019) | 50.781.17 | 49.641.49 | 49.891.67 | 47.940.69 | 69.171.58 | 51.731.94 | |
| Transformer Vaswani et al. (2017) | 50.472.14 | 49.131.83 | 48.011.53 | 48.091.59 | 67.931.59 | 48.932.02 | |
| Medformer Wang et al. (2024) | 53.271.54 | 51.021.57 | 50.711.55 | 50.651.51 | 70.931.19 | 51.211.32 | |
| MedGNN Fan et al. (2025) | 56.120.11 | 55.070.09 | 55.470.34 | 55.000.24 | 74.680.33 | - | |
| HM-BiTCN + CIF | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 | |
| PTB (2-Classes) Reported | Autoformer Wu et al. (2021) | 73.352.10 | 72.112.89 | 63.243.17 | 63.693.84 | 78.543.48 | 74.253.53 |
| Crossformer Zhang and Yan (2022) | 80.173.79 | 85.041.83 | 71.256.29 | 72.757.19 | 88.553.45 | 87.313.25 | |
| FEDformer Zhou et al. (2022) | 76.052.54 | 77.583.61 | 66.103.55 | 67.144.37 | 85.934.31 | 82.595.42 | |
| Informer Zhou et al. (2021) | 78.691.68 | 82.871.02 | 69.192.90 | 70.843.47 | 92.090.53 | 90.020.60 | |
| iTransformer Liu et al. (2024b) | 83.890.71 | 88.251.18 | 76.391.01 | 79.061.06 | 91.181.16 | 90.930.98 | |
| MTST Zhang et al. (2024) | 76.591.90 | 79.881.90 | 66.312.95 | 67.383.71 | 86.862.75 | 83.752.84 | |
| Nonformer Liu et al. (2022) | 78.660.49 | 82.770.86 | 69.120.87 | 70.901.00 | 89.372.51 | 86.672.38 | |
| PatchTST Nie et al. (2023) | 74.741.62 | 76.941.51 | 63.892.71 | 64.363.38 | 88.790.91 | 83.390.96 | |
| Reformer Kitaev et al. (2019) | 77.962.13 | 81.721.61 | 68.203.35 | 69.653.88 | 91.130.74 | 88.421.30 | |
| Transformer Vaswani et al. (2017) | 77.371.02 | 81.840.66 | 67.141.80 | 68.472.19 | 90.081.76 | 87.221.68 | |
| Medformer Wang et al. (2024) | 83.502.01 | 85.190.94 | 77.113.39 | 79.183.31 | 92.811.48 | 90.321.54 | |
| MedGNN Fan et al. (2025) | 84.530.28 | 87.350.45 | 77.900.66 | 80.400.62 | 93.310.46 | - | |
| HM-BiTCN + CIF | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 | |
| PTB-XL (5-Classes) Reported | Autoformer Wu et al. (2021) | 61.682.72 | 51.601.64 | 49.101.52 | 48.852.27 | 82.041.44 | 51.931.71 |
| Crossformer Zhang and Yan (2022) | 73.300.14 | 65.060.35 | 61.230.33 | 62.590.14 | 90.020.06 | 67.430.22 | |
| FEDformer Zhou et al. (2022) | 57.209.47 | 52.386.09 | 49.047.26 | 47.898.44 | 82.134.17 | 52.317.03 | |
| Informer Zhou et al. (2021) | 71.430.32 | 62.640.60 | 59.120.47 | 60.440.43 | 88.650.09 | 64.760.17 | |
| iTransformer Liu et al. (2024b) | 69.280.22 | 59.590.45 | 54.620.18 | 56.200.19 | 86.710.10 | 60.270.21 | |
| MTST Zhang et al. (2024) | 72.140.27 | 63.840.72 | 60.010.81 | 61.430.38 | 88.970.33 | 65.830.51 | |
| Nonformer Liu et al. (2022) | 70.560.55 | 61.570.66 | 57.750.72 | 59.100.66 | 88.320.36 | 63.400.79 | |
| PatchTST Nie et al. (2023) | 73.230.25 | 65.700.64 | 60.820.76 | 62.610.34 | 89.740.19 | 67.320.22 | |
| Reformer Kitaev et al. (2019) | 71.720.43 | 63.121.02 | 59.200.75 | 60.690.18 | 88.800.24 | 64.720.47 | |
| Transformer Vaswani et al. (2017) | 70.590.44 | 61.570.65 | 57.620.35 | 59.050.25 | 88.210.16 | 63.360.29 | |
| Medformer Wang et al. (2024) | 72.870.23 | 64.140.42 | 60.600.46 | 62.020.37 | 89.660.13 | 66.390.22 | |
| MedGNN Fan et al. (2025) | 73.870.18 | 66.260.29 | 61.130.29 | 62.540.20 | 90.210.15 | - | |

### v2

@@ 4.1 Results of Subject-Dependent
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC | ||||
Autoformer Wu et al. (2021) | 87.831.62 | 87.631.66 | 87.221.97 | 87.381.79 | 96.590.88 | 93.821.64 | ||||
| Crossformer Zhang & Yan (2022) | 89.351.32 | 89.001.44 | 88.791.37 | 88.881.40 | 97.520.58 | 95.451.03 | |||||
| FEDformer Zhou et al. (2022) | 77.632.37 | 76.762.17 | 76.682.48 | 76.602.46 | 91.671.34 | 84.942.11 | |||||
| Informer Zhou et al. (2021) | 90.930.90 | 90.740.71 | 90.501.14 | 90.600.94 | 98.190.27 | 96.510.49 | |||||
| iTransformer Liu et al. (2024b) | 64.900.25 | 62.530.27 | 62.210.26 | 62.250.33 | 81.520.29 | 68.870.49 | |||||
| MTST Zhang et al. (2024) | 65.080.69 | 63.850.80 | 62.710.64 | 63.030.58 | 81.360.56 | 69.340.89 | |||||
| Nonformer Liu et al. (2022) | 96.120.47 | 95.940.56 | 95.990.38 | 95.960.47 | 99.590.09 | 99.080.16 | |||||
| PatchTST Nie et al. (2023) | 66.260.40 | 65.080.41 | 64.970.51 | 64.950.42 | 83.070.45 | 71.700.61 | |||||
| Reformer Kitaev et al. (2019) | 91.511.75 | 91.151.79 | 91.651.56 | 91.141.83 | 98.850.35 | 97.880.60 | |||||
| Transformer Vaswani et al. (2017) | 97.000.43 | 96.870.53 | 96.860.36 | 96.860.44 | 99.750.04 | 99.420.07 | |||||
| Medformer Wang et al. (2024a) | 97.620.34 | 97.530.33 | 97.480.40 | 97.500.36 | 99.830.05 | 99.620.12 | |||||
| MedGNN Fan et al. (2025) | 98.420.04 | 98.310.02 | 98.290.05 | 98.300.12 | 99.930.11 | - | |||||
| Medformer + CIF | 98.870.26 | 98.770.27 | 98.860.27 | 98.810.27 | 99.960.01 | 99.920.03 | |||||
@@
We reproduced 12 baselines. Table 1 lists their reported results. Experimental results show that integrating the CIF method into MedGNN and Medformer outperforms existing approaches, fully demonstrating its effectiveness and superiority.
@@ 4.2 Results of Subject-Independent
| Datasets | Models | Accuracy | Precision | Recall | F1 score | AUROC | AUPRC | ||||
Autoformer Wu et al. (2021) | 68.641.82 | 68.482.10 | 68.772.27 | 68.061.94 | 75.943.61 | 74.384.05 | ||||
| Crossformer Zhang & Yan (2022) | 73.771.95 | 79.294.36 | 68.861.70 | 68.931.85 | 72.393.33 | 72.053.65 | |||||
| FEDformer Zhou et al. (2022) | 74.942.15 | 74.591.50 | 73.563.55 | 73.513.39 | 83.721.97 | 82.942.37 | |||||
| Informer Zhou et al. (2021) | 73.114.40 | 75.176.06 | 69.174.56 | 69.475.06 | 70.464.91 | 70.755.27 | |||||
| iTransformer Liu et al. (2024b) | 74.551.66 | 74.772.10 | 71.761.72 | 72.301.79 | 85.591.55 | 84.391.57 | |||||
| MTST Zhang et al. (2024) | 71.141.59 | 79.300.97 | 65.272.28 | 64.013.16 | 68.872.34 | 71.061.60 | |||||
| Nonformer Liu et al. (2022) | 71.893.81 | 71.804.58 | 69.443.56 | 69.743.84 | 70.552.96 | 70.784.08 | |||||
| PatchTST Nie et al. (2023) | 67.031.65 | 78.761.28 | 59.912.02 | 55.973.10 | 65.650.28 | 67.990.76 | |||||
| Reformer Kitaev et al. (2019) | 78.702.00 | 82.503.95 | 75.001.61 | 75.931.82 | 73.941.40 | 76.041.14 | |||||
| Transformer Vaswani et al. (2017) | 76.304.72 | 77.645.95 | 73.095.01 | 73.755.38 | 72.506.60 | 73.237.60 | |||||
| Medformer Wang et al. (2024a) | 78.740.64 | 81.110.84 | 75.400.66 | 76.310.71 | 83.200.91 | 83.660.92 | |||||
| MedGNN Fan et al. (2025) | 82.600.35 | 87.700.22 | 78.930.09 | 80.250.16 | 85.930.26 | - | |||||
| HM-BiTCN + CIF | 85.161.55 | 84.761.62 | 85.331.27 | 84.821.49 | 94.061.07 | 94.210.99 | |||||
Autoformer Wu et al. (2021) | 87.333.79 | 88.063.56 | 87.333.79 | 87.263.84 | 93.812.26 | 93.322.42 | ||||
| Crossformer Zhang & Yan (2022) | 81.562.19 | 81.972.25 | 81.562.19 | 81.502.20 | 91.201.78 | 91.511.71 | |||||
| FEDformer Zhou et al. (2022) | 78.131.98 | 78.521.91 | 78.131.98 | 78.042.01 | 86.561.86 | 86.481.99 | |||||
| Informer Zhou et al. (2021) | 89.022.50 | 89.432.14 | 89.022.50 | 88.982.54 | 96.640.68 | 96.750.63 | |||||
| iTransformer Liu et al. (2024b) | 74.671.06 | 74.711.06 | 74.671.06 | 74.651.06 | 83.371.14 | 83.731.27 | |||||
| MTST Zhang et al. (2024) | 76.963.76 | 77.243.59 | 76.963.76 | 76.883.83 | 85.274.46 | 82.815.64 | |||||
| Nonformer Liu et al. (2022) | 87.882.48 | 88.861.84 | 87.882.48 | 87.782.56 | 97.050.68 | 96.990.68 | |||||
| PatchTST Nie et al. (2023) | 79.253.79 | 79.604.09 | 79.253.79 | 79.203.77 | 87.954.96 | 86.366.67 | |||||
| Reformer Kitaev et al. (2019) | 87.922.01 | 88.641.40 | 87.922.01 | 87.852.08 | 96.300.54 | 96.400.45 | |||||
| Transformer Vaswani et al. (2017) | 87.171.67 | 87.991.68 | 87.171.67 | 87.101.68 | 96.280.92 | 96.340.81 | |||||
| Medformer Wang et al. (2024a) | 89.620.81 | 89.680.78 | 89.620.81 | 89.620.81 | 96.410.35 | 96.510.33 | |||||
| MedGNN Fan et al. (2025) | 91.040.09 | 91.150.12 | 91.040.20 | 91.040.08 | 96.740.04 | - | |||||
| HM-BiTCN + CIF | 93.131.41 | 93.331.37 | 93.131.41 | 93.121.42 | 98.620.66 | 98.680.63 | |||||
Autoformer Wu et al. (2021) | 45.251.48 | 43.671.94 | 42.962.03 | 42.591.85 | 61.021.82 | 43.102.30 | ||||
| Crossformer Zhang & Yan (2022) | 50.452.31 | 45.571.63 | 45.881.82 | 45.501.70 | 66.452.03 | 48.332.05 | |||||
| FEDformer Zhou et al. (2022) | 46.300.59 | 46.050.76 | 44.221.38 | 43.911.37 | 62.621.75 | 46.111.44 | |||||
| Informer Zhou et al. (2021) | 48.451.96 | 46.541.68 | 46.061.84 | 45.741.38 | 65.871.27 | 47.601.30 | |||||
| iTransformer Liu et al. (2024b) | 52.601.59 | 46.791.27 | 47.281.29 | 46.791.13 | 67.261.16 | 49.531.21 | |||||
| MTST Zhang et al. (2024) | 45.602.03 | 44.701.33 | 45.051.30 | 44.311.74 | 62.500.81 | 45.160.85 | |||||
| Nonformer Liu et al. (2022) | 49.951.05 | 47.710.97 | 47.461.50 | 46.961.35 | 66.231.37 | 47.331.78 | |||||
| PatchTST Nie et al. (2023) | 44.370.95 | 42.401.13 | 42.061.48 | 41.971.37 | 60.081.50 | 42.491.79 | |||||
| Reformer Kitaev et al. (2019) | 50.781.17 | 49.641.49 | 49.891.67 | 47.940.69 | 69.171.58 | 51.731.94 | |||||
| Transformer Vaswani et al. (2017) | 50.472.14 | 49.131.83 | 48.011.53 | 48.091.59 | 67.931.59 | 48.932.02 | |||||
| Medformer Wang et al. (2024a) | 53.271.54 | 51.021.57 | 50.711.55 | 50.651.51 | 70.931.19 | 51.211.32 | |||||
| MedGNN Fan et al. (2025) | 56.120.11 | 55.070.09 | 55.470.34 | 55.000.24 | 74.680.33 | - | |||||
| HM-BiTCN + CIF | 58.560.93 | 55.650.81 | 55.860.79 | 55.420.82 | 76.070.59 | 59.750.67 | |||||
Autoformer Wu et al. (2021) | 73.352.10 | 72.112.89 | 63.243.17 | 63.693.84 | 78.543.48 | 74.253.53 | ||||
| Crossformer Zhang & Yan (2022) | 80.173.79 | 85.041.83 | 71.256.29 | 72.757.19 | 88.553.45 | 87.313.25 | |||||
| FEDformer Zhou et al. (2022) | 76.052.54 | 77.583.61 | 66.103.55 | 67.144.37 | 85.934.31 | 82.595.42 | |||||
| Informer Zhou et al. (2021) | 78.691.68 | 82.871.02 | 69.192.90 | 70.843.47 | 92.090.53 | 90.020.60 | |||||
| iTransformer Liu et al. (2024b) | 83.890.71 | 88.251.18 | 76.391.01 | 79.061.06 | 91.181.16 | 90.930.98 | |||||
| MTST Zhang et al. (2024) | 76.591.90 | 79.881.90 | 66.312.95 | 67.383.71 | 86.862.75 | 83.752.84 | |||||
| Nonformer Liu et al. (2022) | 78.660.49 | 82.770.86 | 69.120.87 | 70.901.00 | 89.372.51 | 86.672.38 | |||||
| PatchTST Nie et al. (2023) | 74.741.62 | 76.941.51 | 63.892.71 | 64.363.38 | 88.790.91 | 83.390.96 | |||||
| Reformer Kitaev et al. (2019) | 77.962.13 | 81.721.61 | 68.203.35 | 69.653.88 | 91.130.74 | 88.421.30 | |||||
| Transformer Vaswani et al. (2017) | 77.371.02 | 81.840.66 | 67.141.80 | 68.472.19 | 90.081.76 | 87.221.68 | |||||
| Medformer Wang et al. (2024a) | 83.502.01 | 85.190.94 | 77.113.39 | 79.183.31 | 92.811.48 | 90.321.54 | |||||
| MedGNN Fan et al. (2025) | 84.530.28 | 87.350.45 | 77.900.66 | 80.400.62 | 93.310.46 | - | |||||
| HM-BiTCN + CIF | 88.291.45 | 90.661.48 | 83.212.02 | 85.591.96 | 94.280.93 | 93.781.11 | |||||
Autoformer Wu et al. (2021) | 61.682.72 | 51.601.64 | 49.101.52 | 48.852.27 | 82.041.44 | 51.931.71 | ||||
| Crossformer Zhang & Yan (2022) | 73.300.14 | 65.060.35 | 61.230.33 | 62.590.14 | 90.020.06 | 67.430.22 | |||||
| FEDformer Zhou et al. (2022) | 57.209.47 | 52.386.09 | 49.047.26 | 47.898.44 | 82.134.17 | 52.317.03 | |||||
| Informer Zhou et al. (2021) | 71.430.32 | 62.640.60 | 59.120.47 | 60.440.43 | 88.650.09 | 64.760.17 | |||||
| iTransformer Liu et al. (2024b) | 69.280.22 | 59.590.45 | 54.620.18 | 56.200.19 | 86.710.10 | 60.270.21 | |||||
| MTST Zhang et al. (2024) | 72.140.27 | 63.840.72 | 60.010.81 | 61.430.38 | 88.970.33 | 65.830.51 | |||||
| Nonformer Liu et al. (2022) | 70.560.55 | 61.570.66 | 57.750.72 | 59.100.66 | 88.320.36 | 63.400.79 | |||||
| PatchTST Nie et al. (2023) | 73.230.25 | 65.700.64 | 60.820.76 | 62.610.34 | 89.740.19 | 67.320.22 | |||||
| Reformer Kitaev et al. (2019) | 71.720.43 | 63.121.02 | 59.200.75 | 60.690.18 | 88.800.24 | 64.720.47 | |||||
| Transformer Vaswani et al. (2017) | 70.590.44 | 61.570.65 | 57.620.35 | 59.050.25 | 88.210.16 | 63.360.29 | |||||
| Medformer Wang et al. (2024a) | 72.870.23 | 64.140.42 | 60.600.46 | 62.020.37 | 89.660.13 | 66.390.22 | |||||
| MedGNN Fan et al. (2025) | 73.870.18 | 66.260.29 | 61.130.29 | 62.540.20 | 90.210.15 | - | |||||
@@
Table 2 presents the results reported by various methods in the subject-independent setting. Our method achieves the highest average scores across six metrics on four out of the five datasets. On PTB-XL, our method tops AUROC and AUPRC and ranks second in Accuracy versus reported results, and ranks first in Accuracy, Precision, AUROC, AUPRC, and second in Recall versus our reproduced results. Additionally, it is worth noting that in the subject-independent setup, the F1 score of ADFTD is 55.42%, which is significantly lower than the 99.59% achieved in the subject-dependent setup. This comparison highlights the challenges of the subject-independent setup, which better simulates real-world scenarios.